OpenStudy (anonymous):

Solve 3|2-x|≥12 x ≤ -2 or x ≥ 6 x ≥ 2 and x ≤ 6 x ≤ -2 x ≥ 6

OpenStudy (anonymous):

er???? what do you need ?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

i think its x</= -2

OpenStudy (anonymous):

how do you got 2 and 6? @.@

OpenStudy (lynfran):

\[3|2-x|\ge 12\]\[|2-x|\ge 12/3\]\[|2-x|\ge4\]

OpenStudy (anonymous):

-x>/= 2 divide by -1 x</= -2

OpenStudy (lynfran):

\[|2-x|\ge 4\]\[2-x \ge \pm 4\]\[2-x \ge 4\]and \[2-x \ge -4\]

OpenStudy (lynfran):

\[-x \ge 4-2\]\[-x \ge 2\]\[x \le 2\] and \[-x \ge -4-2\]\[-x \ge -6\]\[x \le 6\]

OpenStudy (lynfran):

error shd be \[x \le-2\]